A)Draw a scatter plot to show a moderately strong relationship between wind consumption and heart disease death rate.
B)Calculate a, b, and R2 in the relationship Y = a + bX, where X is wine consumption and Y is heart disease death rate.
C)Use the equation above to predict the heart disease death rate in a country where wine consumption is 1 and and in a country where the consumption is 8.
Table 2; Problem 6.5 Country Wine Heart diseaseCountryWine Heart disease 7.50 Netherlands 2.75 New Zealand 2.50 Norway1.75 5.00 2.50 Switzerland 6.00 United Kingdom 2.75 United States 1.25 60 70 100 80 50 90 70 120 120 Italy Australia 3.25 Austria 4.75 Belgium 2.75 Canada 1.50 Denmark 4.50 Finland 3.00 France 8.50 Germany 3.75 Iceland1.25 Ireland 2.00 80 100 60 80 90 120 40 90 110 130 Spain Sweden
Answer #1

a)

Heart disease 140 120 100 80 60 Heart disease 40 20 0.00 2.00 4.00 6.00 8.00 10.00

b) solving using excel

SUMMARY OUTPUT Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.644904211 0.415901441

a = 115.86

b = 8.05

R Square =0.416

\small \hat{Y}= 115.86 - 8.05X

c) Predicted heart disease death rate in a country where wine consumption is 1

\small \hat{Y} = 115.86 - 8.05(1)

= 107.81

Predicted heart disease death rate in a country where the consumption is 8

\small \hat{Y}= 115.86 - 8.05(8)

= 51.46